A particular poll tracks daily the percentage of Americans who approve or disapprove of the performance by President 1. Daily results are based on random telephone interviews with approximately 1100 national adults. The poll reports that 51​% of adults approve of President 1. The same poll reported an approval rating of 52​% for President 2. A news anchor remarks that​ "President 1​ doesn't even get as much approval as President 2​ did." Is there evidence that this difference is​ real? State and test the appropriate hypotheses. Check conditions.

What are the null and alternative​ hypotheses?

Now check the necessary assumptions and conditions.

Test Statistic?

P-Value?
Answer #1

Ans:

H_0:p_1=p_2

H_a:p_1\neq p_2

Test statistic:

z=(0.51-0.52)/sqrt(0.515*(1-0.515)*((1/1100)+(1/1100)))

z=-0.469

p-value(2 tailed)=2*P(z<-0.469)=0.6390

Fail to reject the null hypothesis

There is not sufficient evidence to indicate a significant difference.